# On subgroups with narrow Schreier graphs

### v2

@@ 5 Reducing a Euclidean cube complex to a line
If, in addition to the conditions of the proposition, one assumes for example that $\Gamma$ acts on $Y$ without a fixed point in the visual boundary then [CS11, Theorem E] implies that $\Gamma$ fixes a Euclidean flat in $Y$. The task of finding an appropriate map to $\mathbb{Z}$ then requires some work (and probably some CAT(0) geometry). It is moreover possible to reduce to the case where either there is no fixed point in the visual boundary or there is a fixed point in the Roller boundary [CFI16, Proposition 2.26]. In the later case, Caprace’s appendix in [CFI16] provides several epimorphisms from a finite index subgroup of $\Gamma$ to $\mathbb{Z}$ (in fact, in either case one can show that $\Gamma$ virtually fixes the simplicial boundary of $Y$ pointwise which suffices to apply Caprace’s argument). The fact that one of these has a kernel with the required properties is not immediate, but it can be shown following much the same argument as the proof of Proposition 1.11.
@@ 6 A family of examples
A family of groups introduced by Houghton [Hou78] provides an example, for each $n\geq 2$, of a finitely generated pair of groups $H_{n}\leq G_{n}$ such that $H_{n}\operatorname{\setminus}G_{n}$ is an $n$-ended quasi-tree with linear growth – and therefore narrow by Proposition 3.2 – but $H_{n}$ is not a virtual fiber subgroup. These groups are known to provide examples of pairs of groups with an arbitrary number of relative ends [Sco77, Example 2.1] and, as will become apparent, the fact that $H_{n}\operatorname{\setminus}G_{n}$ is not only $n$-ended but a quasi-tree with linear growth is obvious from the construction.
@@
Fix $n\geq 3$ and let $\sigma$ be a permutation of $\{1,\dots,n\}$. Define $\alpha:X_{n}\rightarrow X_{n}$ by $\alpha(i,m)=(\sigma(i),m)$ and let $L$ be the subgroup of the permutation group of $X_{n}$ generated by $S_{n}\cup\{\alpha\}$. Note that $\alpha\mathcal{G}_{n}\alpha^{-1}=\mathcal{G}_{n}$ and $\alpha\notin\mathcal{G}_{n}$ unless $\sigma=\operatorname{id}$. Therefore $L=\mathcal{G}_{n}\rtimes\mathbb{Z}/m\mathbb{Z}$, where $m$ is the order of $\sigma$, and $L$ is of type $F_{n-1}$ by [Alo94, Corollary 9]. Let $M\coloneqq\operatorname{Stab}_{L}((1,1))$. Then $\alpha\in M$ so $M=H_{n}\rtimes\operatorname{\langle}\alpha\operatorname{\rangle}\cong\mathcal{G}_{n}\rtimes\mathbb{Z}/m\mathbb{Z}=L$. Moreover, the Schreier graph of $M$ with respect to $S_{n}\cup\{\alpha\}$ is a quasi-tree whose number of ends is the number of orbits of the action of $\operatorname{\langle}\sigma\operatorname{\rangle}$ on $\{1,\dots,n\}$ and $L=M\cup Mg_{1}M$.
Using cyclic notation, if we take for instance $\sigma=(2,\dots,n)$ then this provides an example of a pair $H\leq G$ where $H\operatorname{\setminus}G$ is a quasi-line, $H$ is not a virtual fiber subgroup and both $H$ and $G$ are of type $F_{n-1}$.

### v3

@@ 5 Reducing a Euclidean cube complex to a line
If, in addition to the conditions of the proposition, one assumes for example that $\Gamma$ acts on $Y$ without a fixed point in the visual boundary then [CS11, Theorem E] implies that $\Gamma$ fixes a Euclidean flat in $Y$. The task of finding an appropriate map to $\mathbb{Z}$ then requires some work (and probably some CAT(0) geometry). It is moreover possible to reduce to the case where either there is no fixed point in the visual boundary or there is a fixed point in the Roller boundary [CFI16, Proposition 2.26]. In the later case, Caprace’s appendix in [CFI16] provides a number of epimorphisms from a finite index subgroup of $\Gamma$ to $\mathbb{Z}$ (in fact, in either case one can show that $\Gamma$ virtually fixes the simplicial boundary of $Y$ pointwise which suffices to apply Caprace’s argument). The fact that one of these has a kernel with the required properties is not immediate, but it can be shown following much the same argument as the proof of Proposition 1.12.
@@ 6 A family of examples
A family of groups introduced by Houghton [Hou78] provides an example, for each $n\geq 2$, of a finitely generated pair of groups $H_{n}\leq G_{n}$ such that $H_{n}\operatorname{\setminus}G_{n}$ is an $n$-ended quasi-tree with linear growth – and is therefore narrow by Proposition 3.2 – but $H_{n}$ is not a virtual fiber subgroup. These groups are known to provide examples of pairs of groups with an arbitrary number of relative ends [Sco77, Example 2.1] and, as will become apparent, the fact that $H_{n}\operatorname{\setminus}G_{n}$ is not only $n$-ended but a quasi-tree with linear growth is obvious from the construction.
@@
Fix $n\geq 3$ and let $\sigma$ be a permutation of $\{1,\dots,n\}$ such that $\sigma(1)=1$. Define $\alpha:X_{n}\rightarrow X_{n}$ by $\alpha(i,m)=(\sigma(i),m)$ and let $L$ be the subgroup of the permutation group of $X_{n}$ generated by $S_{n}\cup\{\alpha\}$. Note that $\alpha\mathcal{G}_{n}\alpha^{-1}=\mathcal{G}_{n}$ and $\alpha\notin\mathcal{G}_{n}$ unless $\sigma=\operatorname{id}$. Therefore $L=\mathcal{G}_{n}\rtimes\mathbb{Z}/m\mathbb{Z}$, where $m$ is the order of $\sigma$, and $L$ is of type $F_{n-1}$ by [Alo94, Corollary 9]. Let $M\coloneqq\operatorname{Stab}_{L}((1,1))$. Then $\alpha\in M$ so $M=H_{n}\rtimes\operatorname{\langle}\alpha\operatorname{\rangle}$ and $M$ also has type $F_{n-1}$. [^1] In fact, $M\cong L$: Let $\psi:X_{n}\rightarrow X_{n}-\{(1,1)\}$ be the bijection defined by $\psi(i,n)=(i,n)$ if $i\neq 1$ and $\psi(i,n)=(i,n+1)$ otherwise. Define $\overline{\psi}:\mathcal{G}_{n}\rightarrow H_{n}$ as follows. If $f\in\mathcal{G}_{n}$ then $\overline{\psi}(f)((1,1))=(1,1)$ and $\overline{\psi}(f)_{|X_{n}-\{(1,1)\}}=\psi\circ f\circ\psi^{-1}$. One can check that $\overline{\psi}$ is an $\operatorname{\langle}\alpha\operatorname{\rangle}$-equivariant isomorphism so it induces an isomorphism $L\rightarrow M$.
Moreover, the Schreier graph of $M$ with respect to $S_{n}\cup\{\alpha\}$ is a quasi-tree whose number of ends is the number of orbits of the action of $\operatorname{\langle}\sigma\operatorname{\rangle}$ on $\{1,\dots,n\}$ and $L=M\cup Mg_{1}M$. By Theorem 1.1, this implies that $M$ is not a virtual fiber subgroup.
Using cyclic notation, if we take for instance $\sigma=(2,\dots,n)$ then this provides an example of a pair $H\leq G$ where $H\operatorname{\setminus}G$ is a quasi-line, both $H$ and $G$ are of type $F_{n-1}$ and $H$ is not a virtual fiber subgroup.